# A Full AGN Feedback Prescription for Numerical Models: Negative, Positive and Hot Gas-Ejection Modes

### v1

@@ Abstract
We build upon the state-of-the-art semi-analytic model FEGA24 (Formation and Evolution of GAlaxies, Contini et al. 2024b), which integrates the latest prescriptions relevant to galaxy formation and evolution, alongside a comprehensive AGN feedback model. This model incorporates three modes of feedback: negative (preventing excessive cooling), positive (enhancing star formation), and hot gas ejection (expelling gas beyond the virial radius of halos). These modes operate in a coordinated manner: the negative mode regulates the cooling process, the positive mode promotes bursts of star formation, and the hot gas ejection mode expels gas beyond the virial radius when the AGN is sufficiently powerful. Our updated semi-analytic model, FEGA25, retains the qualitative and quantitative consistency of the analyses presented in Contini et al. (2024b), while delivering more robust results. Notably, FEGA25 provides a more detailed characterization of the fraction of red galaxies as a function of stellar mass, predicts a main sequence of star-forming galaxies more consistent with observations, and estimates the fraction of hot gas in halos closer to observed values. These findings underscore the importance of a physical mechanism capable of ejecting hot gas beyond the virialized region of dark matter halos without significantly altering the stellar and cold gas components. Such a mechanism is crucial to ensure the proper functioning of other processes, such as cooling and star formation. Since supernova feedback is already modeled at its maximum efficiency, AGN feedback emerges as the natural candidate for this role.
@@ 1 Introduction
In Section 2, we describe the three AGN feedback modes (radio mode) implemented in the model and provide a brief overview of the calibration process. Section 3 presents a detailed analysis of the key advancements introduced in FEGA25, including refined star formation rate and red fraction–stellar mass relations, as well as a reduction in the predicted hot gas fraction as a function of halo mass. We also explore the impact of the updated positive AGN feedback implementation compared to the previous version. Section 4 discusses the broader implications of our findings, and Section 5 summarizes the main conclusions of this study. Throughout this work, we adopt a Chabrier (Chabrier 2003) initial mass function for stellar mass calculations, and all units are h-corrected.
@@ 2.1 AGN Feedback: Negative, Positive and Ejection Modes
The AGN feedback plays a crucial role in galaxy formation, primarily by regulating the accretion of hot gas onto the BH and controlling the growth of the host galaxy. In the standard framework, AGN feedback operates in two main modes: (1) the quasar mode, responsible for BH growth during galaxy mergers; and (2) the radio mode, which is believed to limit excessive galaxy growth by injecting energy to counteract ongoing cooling. The energy injected is proportional to the accretion of hot gas onto the BH and is governed by the efficiency of this process. For an overview of the quasar mode in FEGA24, we refer the reader to C24, while a more detailed description can be found in the original implementation by Croton et al. (2006).
@@ 2.1.1 Negative Feedback
The injection of mechanical energy described above can entirely halt gas cooling, reducing $\dot{M}_{\rm{cool,new}}$ to zero. However, this mode of AGN feedback operates with varying efficiency over time, being almost negligible at high redshift and becoming increasingly significant at lower redshifts. This shift corresponds to the rapid growth of BHs due to the increasing availability of hot gas, partly resulting from the reincorporation of gas previously ejected beyond the halos’ virial radius. Moreover, the efficiency of this negative mode depends on halo mass, being more effective in halos with $\log M_{\rm{halo}}>13$.
@@ 2.1.2 Positive Feedback
The positive AGN feedback, first introduced in C24, was an innovative attempt to incorporate a physical mechanism into SAMs, supported by various observational studies (Zinn et al. 2013; Salomé et al. 2015; Cresci et al. 2015a, b; Mahoro et al. 2017; Salomé et al. 2017; Shin et al. 2019; Nesvadba et al. 2020; Joseph et al. 2022; Tamhane et al. 2022; Venturi et al. 2023; Gim & Reines 2024) as well as other theoretical and numerical approaches (e.g., Gaibler et al. 2012; Zubovas et al. 2013; Silk 2013; Bieri et al. 2015, 2016; Zubovas & Bourne 2017; Mukherjee et al. 2018; Silk et al. 2024). However, as noted in the previous work, the prescription used was more an attempt to isolate the positive feedback effect rather than a fully physical model. The goal was to demonstrate that both the negative and positive modes of AGN feedback can coexist without altering the overall properties of galaxies and their scaling relations.
@@
Compared to Equation 4, Equation 5 provides a more physically motivated implementation of positive AGN feedback for several reasons. The most significant is that the efficiency of this mode now depends directly on the efficiency of the negative mode through the term $\frac{\delta M_{\rm{BH}}}{M_{\rm{BH}}}\propto\kappa_{\rm{AGN}}$, which incorporates the accretion efficiency onto the BH as governed by $\kappa_{\rm{AGN}}$. This way, the efficiencies of both feedback modes are controlled by a single parameter, $\kappa_{\rm{AGN}}$, enabling the MCMC algorithm to focus on calibrating this parameter. Another important aspect of this approach is that it links the two modes: the positive mode can be strong when the negative mode is weak, and vice versa, or they can both be strong simultaneously. A key expectation is that the positive mode will dominate at high redshifts, when the negative mode is less effective, and will fade at lower redshifts, when the negative mode is more active.
@@ 2.1.3 Hot Gas Ejection
where $V_{\rm{scale}}$ is a free parameter, calibrated during the model fitting process, that controls the efficiency threshold. Larger values of $V_{\rm{scale}}$ enhance the efficiency in larger halos, while smaller values reduce the efficiency in lower-mass halos. The ejected gas mass, $M_{\rm{ejected}}$, is then stored in the ejected reservoir, but it remains available for later reincorporation, as described in C24. Clearly, if $V_{200}>V_{\rm{scale}}$, the amount of hot gas ejected is assumed to be 0.
@@ 2.2.1 Set of Parameters
Building on the previous calibration of FEGA24 ModA from C24 and the subsequent improvements made by the implementation of stellar halos in Contini et al. (2024a), we can complete the final calibration of FEGA25. In this process, we calibrate the relevant parameters within the AGN feedback modes, as well as the parameter $\gamma_{2}$ from C24, which plays a role in the computation of the reincorporation time of ejected gas. This parameter is particularly important due to the introduction of the hot gas ejection mode, which adds more gas to the ejected reservoir. The parameters constrained during the calibration are: the efficiency $\kappa_{\rm{AGN}}$ of the BH in Equation 1, the scale velocity $V_{\rm{scale}}$ in Equation 6, and the reincorporation factor $\gamma_{2}$. The values obtained during calibration are the following for FEGA25 (FEGA24): $\kappa_{\rm{AGN}}=2\rm{e}^{-5}\,(5\rm{e}^{-5})$ $M_{\odot}$/yr; $V_{\rm{scale}}=500$ (not applicable) km/s; $\gamma_{2}=0.378\,(0.375)$.
A quick inspection of the parameter values reveals one key difference between FEGA24 and FEGA25. While the reincorporation factor $\gamma_{2}$ remains nearly unchanged, the efficiency parameter $\kappa_{\rm{AGN}}$ in FEGA25 is 2.5 times smaller than in FEGA24. The minimal change in $\gamma_{2}$ suggests that the continuous addition of hot gas to the ejected reservoir from the hot gas ejection mode does not significantly affect the reincorporation time of the expelled gas. On the other hand, the lower value of $\kappa_{\rm{AGN}}$ means that less hot gas is available for BH accretion, resulting in a lower energy injection required to counterbalance the cooling flow. Consequently, the BH accretes less hot gas. This point will be discussed further in Section 4.
@@ 3 Results
Figure 1: The evolution of the SMF from $z=3$ to $z=0$ is shown, as predicted by FEGA24 (represented by green stars) and FEGA25 (represented by blue squares), compared with the set of observations used to calibrate the models (black lines). Additionally, we compare our predictions with those of the Henriques et al.’s SAM, as well as another set of observations (at $z=0$ only) from different authors: Li & White 2009 (represented by a cyan line), D’Souza et al. 2015 (represented by a purple line), and Bernardi et al. 2018 (represented by magenta lines indicating the ±$\pm\sigma$ region). Both models accurately describe the evolution of the SMF down to $z=0$.
@@
Figure 2: The star formation rate-stellar mass relation at $z=0$ is shown, as predicted by FEGA24 (left panel) and FEGA25 (right panel). Both models are compared with the observed main sequence data from Elbaz et al. 2007 (represented by blue lines). The contours represent different percentages of the data, as indicated by the colors in the accompanying bars. The last contour encloses 90% of the full sample, while the remaining 10% are depicted by individual black dots. Overall, both models describe the observed main sequence well and predict a broad red cloud. However, while FEGA24 is somewhat biased-high compared to the observations, FEGA25 overlaps perfectly with them. The net improvement over the previous version of the model is clear, and this is entirely attributable to our new implementation of AGN feedback.
Figure 3: The fraction of red central galaxies as a function of stellar mass is shown, as predicted by FEGA24 (represented by the green region) and FEGA25 (represented by the blue region), and compared with observed fractions from SDSS-DR7 (De Lucia et al. 2024, represented by magenta and blue diamonds), SDSS-DR8 (Hirschmann et al. 2014, indicated by orange stars), and the GAMA survey (Davies et al. 2019, marked by purple triangles). As in C24, red galaxies are selected based on different cuts in $\log$SSFR, specifically: -10.5 for GAMA data, and -10.6 and -11.0 for SDSS data. The green and blue regions in our models correspond to the smallest and largest SSFR cuts, respectively. This plot highlights a significant improvement over the previous version of the model. FEGA25 not only captures the observed trend in greater detail but, importantly, also predicts the red fraction on dwarf scales. This behavior in FEGA25 can be attributed to the new hot gas ejection mode in the AGN feedback, as explained in the text.
Figure 4: As done in C24, we present the percentage of stellar mass contributed by the positive AGN feedback mode as a function of halo mass (for central galaxies only). The predictions of FEGA24 (left panel) and FEGA25 (right panel) are shown at two redshifts: $z=0$ (green lines) and $z=3$ (black lines). The plots account for all stellar mass attributed to positive AGN feedback by considering all branches of the merger tree for each galaxy. There are notable differences between the positive AGN feedback implemented in C24 (left panel) and the new implementation in FEGA25 (right panel). First, the new positive AGN feedback in FEGA25 naturally and distinctly captures the expected redshift dependence, with higher efficiency at higher redshifts. Secondly, while in C24 there was no clear trend with stellar mass, only a larger scatter at lower mass scales (left panel), FEGA25 shows an increase in the efficiency of positive AGN feedback with stellar mass. This is particularly evident where the negative mode is also efficient, though not strong enough to heat all the cooling gas.
Figure 5: The star formation history of eight sample galaxies, ordered by increasing stellar mass from top to bottom, is shown from redshift $z\sim 10$ to the present time. The red line represents the total SFR, while the blue line indicates the SFR due solely to positive AGN feedback. Several interesting features can be observed in these plots. From top to bottom, it is clear that the contribution of positive AGN feedback to the total SFR increases, as expected from Figure 4. This trend is evident not only in terms of the mass provided but also in the duration of activity of the positive feedback mode. Specifically, in lower-mass galaxies, the contribution is shorter and characterized by episodic bursts, whereas in higher-mass galaxies, the SFR history driven by positive AGN feedback more closely follows the total SFR, being longer and more sustained. In all cases, the positive feedback mode is efficient at high redshift and begins to fade as the negative feedback mode takes over the cooling process, indicating that the negative mode reaches its peak power.
Figure 6: The fraction of mass for different components (stars, cold gas, hot gas, and ejected gas) within the virial radius, normalized by the universal baryon fraction, is shown as a function of halo mass. Different colors represent the various components, and the solid and dashed lines distinguish the predictions from FEGA24 and FEGA25, respectively. Our models are compared with the observed hot gas fraction from Popesso et al. 2024 (magenta diamonds with error bars) and a set of observed hot gas fractions from various authors (Vikhlinin et al. 2006; Arnaud et al. 2007; Sun et al. 2009; Rasmussen & Ponman 2009; Pratt et al. 2009; Mahdavi et al. 2013; Gonzalez et al. 2013; Lovisari et al. 2015; Eckert et al. 2016; Pearson et al. 2017; Eckert et al. 2019; Mulroy et al. 2019; Ragagnin et al. 2022, dark blue circles). It is important to note that, despite the large scatter in the observed data, FEGA24 fails to predict a reasonable hot gas fraction for halos smaller than large groups ($\log M_{\rm{halo}}<14$). In contrast, FEGA25, with its new hot gas ejection mode in AGN feedback, performs significantly better. This result underscores the importance of having a mechanism, such as the hot gas ejection mode, that can expel gas from the halo without significantly altering the cold gas and stellar mass content.
@@
We now proceed with the analysis by examining Figure 1, which depicts the evolution of the SMF from $z=3$ to the present day. The predictions from FEGA25 (blue squares) are compared to those of FEGA24 (green stars), the results from the Henriques et al. (2020) SAM (brown dashed lines), and various observational datasets, as indicated in the legend. As in C24, we remind the reader that the evolution of the SMF was used to calibrate the model. Therefore, these predictions should not be considered fully independent.
@@
The minor differences between the two models can be attributed to several factors. Although the BH accretion efficiency, $\kappa_{\rm{AGN}}$, is 2.5 times lower in FEGA25—which would suggest an opposite trend—the differing implementations of positive AGN feedback in the models, and particularly the new prescription for hot gas ejection in FEGA25, play a crucial role. Due to the ejection of some hot gas driven by AGN feedback, the amount of hot gas available for cooling in large groups and clusters, where AGN feedback is active, may be somewhat reduced. Consequently, $\kappa_{\rm{AGN}}$ could also decrease. The net effect is a lower overall efficiency, as the reduced availability of hot gas is sufficient to suppress cooling flows, which are also expected to be, on average, slightly diminished. We will revisit this topic in Section 4, where a broader perspective will be provided.

### v2

@@ Abstract
We build upon the state-of-the-art semi-analytic model FEGA24 (Formation and Evolution of GAlaxies, Contini et al. 2024b), which integrates the latest prescriptions relevant to galaxy formation and evolution, alongside a comprehensive AGN feedback model. This model incorporates three modes of feedback: negative (preventing excessive cooling), positive (enhancing star formation), and hot gas ejection (expelling gas beyond the virial radius of halos). These modes operate in a coordinated manner: the negative mode regulates the cooling process, the positive mode promotes bursts of star formation, and the hot gas ejection mode expels gas beyond the virial radius when the AGN is sufficiently powerful. Our updated semi-analytic model, FEGA25, retains the qualitative and quantitative consistency of the analyses presented in Contini et al. (2024b), while delivering more robust results. Notably, FEGA25 provides a more detailed characterization of the fraction of passive galaxies as a function of stellar mass and redshift, predicts a main sequence of star-forming galaxies more consistent with observations and a more accurate cosmic star formation rate density with redshift. Moreover, it estimates the fraction of hot gas in halos closer to observed values. These findings underscore the importance of a physical mechanism capable of ejecting hot gas beyond the virialized region of dark matter halos without significantly altering the stellar and cold gas components. Such a mechanism is crucial to ensure the proper functioning of other processes, such as cooling and star formation. Since supernova feedback is already modeled at its maximum efficiency, AGN feedback emerges as the natural candidate for this role.
@@ 1 Introduction
In Section 2, we describe the three AGN feedback modes (radio mode) implemented in the model and provide a brief overview of the calibration process. Section 3 presents a detailed analysis of the key advancements introduced in FEGA25, including refined star formation rate and passive fraction–stellar mass relations, as well as a reduction in the predicted hot gas fraction as a function of halo mass. We also explore the impact of the updated positive AGN feedback implementation compared to the previous version. This section discusses also the broader implications of our findings, while Section 4 summarizes the main conclusions of this study. Throughout this work, we adopt a Chabrier (Chabrier 2003) initial mass function for stellar mass calculations, and all units are h-corrected.
@@ 2.1 AGN Feedback: Negative, Positive and Ejection Modes
The AGN feedback plays a crucial role in galaxy formation, primarily by regulating the accretion of hot gas onto the BH and controlling the growth of the host galaxy. In the standard framework, AGN feedback operates in two main modes: (1) the quasar mode, responsible for BH growth during galaxy mergers; and (2) the radio mode, which is believed to limit excessive galaxy growth by injecting energy to counteract ongoing cooling. The energy injected is proportional to the accretion of hot gas onto the BH and is governed by the efficiency of this process. For an overview of the quasar mode in FEGA24, we refer the reader to C24, while a more detailed description can be found in the original implementation by Croton et al. (2006). Here, it is worth mentioning that BHs naturally form during the quasar mode through the accretion of cold gas, in an amount defined by Equation 20 in C24. They subsequently grow through both the radio mode (hot gas accretion) and the quasar mode (cold gas accretion and mergers).
@@ 2.1.1 Negative Feedback
The injection of mechanical energy described above can entirely halt gas cooling, by reducing $\dot{M}_{\rm{cool,new}}$ to zero. However, this mode of AGN feedback operates with varying efficiency over time, being almost negligible at high redshifts and becoming increasingly significant at lower redshifts, favored by the increasing availability of hot gas, partly resulting from the reincorporation of gas previously ejected beyond the halos’ virial radius. Moreover, the efficiency of this negative mode depends on halo mass, being more effective in halos with $\log M_{\rm{halo}}>13$.
@@ 2.1.2 Positive Feedback
The positive AGN feedback, first introduced in C24, was an innovative attempt to incorporate a physical mechanism into SAMs, supported by various observational studies (Zinn et al. 2013; Salomé et al. 2015; Cresci et al. 2015a, b; Mahoro et al. 2017; Salomé et al. 2017; Shin et al. 2019; Nesvadba et al. 2020; Joseph et al. 2022; Tamhane et al. 2022; Venturi et al. 2023; Gim & Reines 2024) as well as other theoretical and numerical approaches (e.g., Gaibler et al. 2012; Zubovas et al. 2013; Silk 2013; Bieri et al. 2015, 2016; Zubovas & Bourne 2017; Mukherjee et al. 2018; Silk et al. 2024), where the AGN can trigger star formation by induced pressure at the edge of the outflows (Cresci et al. 2015a), or enhance star formation via burst in different locations depending on the power of the feedback (e.g., Silk 2013). However, as noted in the previous work, the prescription used was more an attempt to isolate the positive feedback effect rather than a fully physical model. The goal was to demonstrate that both the negative and positive modes of AGN feedback can coexist without altering the overall properties of galaxies and their scaling relations.
@@
Compared to Equation 4, Equation 5 provides a more physically-motivated implementation of positive AGN feedback for several reasons. The most significant is that the efficiency of this mode now depends directly on the efficiency of the negative mode through the term $\frac{\delta M_{\rm{BH}}}{M_{\rm{BH}}}\propto\kappa_{\rm{AGN}}$, which incorporates the accretion efficiency onto the BH as governed by $\kappa_{\rm{AGN}}$. This way, the efficiencies of both feedback modes are controlled by a single parameter, $\kappa_{\rm{AGN}}$, enabling the MCMC algorithm to focus on calibrating this parameter. Another important aspect of this approach is that it links the two modes: the positive mode can be strong when the negative mode is weak, and vice versa, or they can both be strong simultaneously. A key expectation is that the positive mode will dominate at high redshifts, when the negative mode is less effective, and will fade at lower redshifts, when the negative mode is more active.
@@ 2.1.3 Hot Gas Ejection
where $V_{\rm{scale}}$ is a free parameter, calibrated during the model fitting process, that controls the efficiency threshold.  [^1] This prescription is applied even when $\dot{M}_{\rm{cool}}>2\frac{\dot{E}_{\rm{radio}}}{V_{200}^{2}}$, i.e., when the injection of energy is not sufficient to completely stop the cooling of gas. Larger values of $V_{\rm{scale}}$ enhance the efficiency in larger halos, while smaller values reduce the efficiency in lower-mass halos. The ejected gas mass, $M_{\rm{ejected}}$, is then stored in the ejected reservoir, but it remains available for later reincorporation, as described in C24. Clearly, if $V_{200}>V_{\rm{scale}}$, the amount of hot gas ejected is assumed to be 0.
@@ 2.2.1 Set of Parameters
Building on the previous calibration of FEGA24 ModA from C24 and the subsequent improvements made by the implementation of stellar halos in Contini et al. (2024a), we can complete the final calibration of FEGA25. In this process, we calibrate the relevant parameters within the AGN feedback modes, as well as the parameter $\gamma_{2}$  [^2] This parameter enters in the definition of the mass of ejected gas that turns back to the hot gas available inside the virial radius, as follows: $\dot{M}_{\rm{ej}}=\gamma_{2}\frac{M_{\rm{ej}}}{t_{\rm{reinc}}}$, where $M_{\rm{ej}}$ is the available gas in the ejecta and $t_{\rm{reinc}}$ is the reicorporation time. More details can be found in C24. from C24, which plays a role in the computation of the reincorporation time of ejected gas. This parameter is particularly important due to the introduction of the hot gas ejection mode, which adds more gas to the ejected reservoir. The parameters constrained during the calibration are: the efficiency $\kappa_{\rm{AGN}}$ of the BH in Equation 1, the scale velocity $V_{\rm{scale}}$ in Equation 6, and the reincorporation factor $\gamma_{2}$. The values obtained during calibration are the following for FEGA25 (FEGA24): $\kappa_{\rm{AGN}}=2\rm{e}^{-5}\,(5\rm{e}^{-5})$ $M_{\odot}$/yr; $V_{\rm{scale}}=500$ (not applicable) km/s; $\gamma_{2}=0.378\,(0.375)$.
A quick inspection of the parameter values reveals one key difference between FEGA24 and FEGA25. While the reincorporation factor $\gamma_{2}$ remains nearly unchanged, the efficiency parameter $\kappa_{\rm{AGN}}$ in FEGA25 is 2.5 times smaller than in FEGA24. The minimal change in $\gamma_{2}$ suggests that the continuous addition of hot gas to the ejected reservoir from the hot gas ejection mode does not significantly affect the reincorporation time of the expelled gas. On the other hand, the lower value of $\kappa_{\rm{AGN}}$ means that less hot gas is available for BH accretion, resulting in a lower energy injection required to counterbalance the cooling flow. Consequently, the BH accretes less hot gas.
@@ 3 Results and Discussion
Figure 1: The evolution of the SMF from $z=3$ to $z=0$ is shown, as predicted by FEGA24 (represented by green stars) and FEGA25 (represented by blue squares), compared with the set of observations used to calibrate the models (black lines). Additionally, we compare our predictions with those of L-Galaxies SAM (Henriques et al. 2020), as well as another set of observations (at $z=0$ only) from different authors: Li & White 2009 (represented by a cyan line), D’Souza et al. 2015 (represented by a purple line), and Bernardi et al. 2018 (represented by magenta lines indicating the ±$\pm\sigma$ region). Both models accurately describe the evolution of the SMF down to $z=0$.
In this section, we present our analysis focused on the key advancements introduced by the new implementations, with particular emphasis on the role of the third AGN feedback mode—namely, the hot gas ejection mode. In Section 3.1, we briefly describe and comment on the evolution of the SMF, which serves as the main observable for calibrating both models. Section 3.2 highlights the primary differences between the models, based on a repetition of the analysis performed in C24. Finally, in Section 3.3, we delve deeper into the specific impact of the positive AGN feedback and the hot gas ejection mode on galaxy evolution.
Given the importance of the topics addressed, it is worth reiterating the fundamental differences between FEGA25 and its predecessor, FEGA24. The first significant update in FEGA25 involves a more detailed description of stellar halos surrounding central galaxies (Contini et al. 2024a). In this framework, these stellar halos form through the reincorporation of stars originally belonging to the diffuse light or intracluster light (ICL) in a transition region between the ICL and the central galaxy. However, this implementation does not influence the analysis conducted in this work.
The second major update is particularly important and represents an unprecedented advancement in semianalytic modeling. While FEGA24 isolated the positive AGN feedback using Equation 4, FEGA25 introduces a more physically-motivated approach, described by Equation 5. This new formulation links the positive AGN feedback to its negative counterpart, both controlled by a single free parameter, the BH accretion efficiency, $\kappa_{\rm{AGN}}$.
3.1 Evolution of the Stellar Mass Function
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Figure 2: The star formation rate-stellar mass relation at $z=0,1,2,3$ is shown, as predicted by FEGA24 (left panels) and FEGA25 (right panels). Both models are compared with the observed main sequence data from Elbaz et al. 2007 (represented by blue lines), Daddi et al. 2007 (blue lines), Karim et al. 2011 (purple diamonds) and Whitaker et al. 2014 (brown tringles). The contours represent different percentages of the data, as indicated by the colors in the accompanying bars. The last contour encloses 90% of the full sample, while the remaining 10% are depicted by individual black dots. Both models are able to catch the overall evolution of this relation, although (apparently, see text for more details) FEGA24 seems to be more in line with observations.
@@
We proceed with the analysis by examining Figure 1, which depicts the evolution of the SMF from $z=3$ to the present day. The predictions from FEGA25 (blue squares) are compared to those of FEGA24 (green stars), the results from L-Galaxies SAM (Henriques et al. 2020, brown dashed lines), and various observational datasets, as indicated in the legend. As in C24, we remind the reader that the evolution of the SMF was used to calibrate the model. Therefore, these predictions should not be considered fully independent.
@@
The minor differences between the two models can be attributed to several factors. Although the BH accretion efficiency, $\kappa_{\rm{AGN}}$, is 2.5 times lower in FEGA25—which would suggest an opposite trend—the differing implementations of positive AGN feedback in the models, and particularly the new prescription for hot gas ejection in FEGA25, play a crucial role. Due to the ejection of some hot gas driven by AGN feedback, the amount of hot gas available for cooling in large groups and clusters, where AGN feedback is active, may be somewhat reduced. Consequently, $\kappa_{\rm{AGN}}$ could also decrease. The net effect is a lower overall efficiency, as the reduced availability of hot gas is sufficient to suppress cooling flows, which are also expected to be, on average, slightly diminished.
That said, one caveat must be noted. As discussed in C24, the high-mass end of the SMF at high redshift is somewhat underpredicted due to the limited box size, which may not be large enough to capture a representative sample of massive galaxies at those redshifts. Additionally, it is important to highlight that our data reflect the raw number densities produced by the models, rather than fitted values.
Figure 3: Cosmic SFR density as a function of redshift as predicted by FEGA24 (green line) and FEGA25 (blue line), compared with the observed data from Driver et al. (2018), (red crosses) and from D’Silva et al. (2023) (magenta diamonds). FEGA25 is more capable to follow the observed data than FEGA24 at redshift $z>2$. This plot accounts for all galaxies in YS200 having $\log M_{*}\geq 8.0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: The fraction of passive galaxies as a function of stellar mass, at $z=0.2,1,2,3$ as predicted by FEGA24 (represented by the green regions) and FEGA25 (represented by blue regions), compared withn the observational data from Muzzin et al. 2013 green regions and from Weaver et al. 2023 (red diamonds). Both model can describe the evolution with redshift of the passive fraction of galaxies up to $z=2$, with FEGA25 being more accurate, but they both underestimate the observed fractions at $z=3$.
3.2 FEGA24 vs FEGA25